OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2-6x+8=y

OpenStudy (anonymous):

First, set it equal to 0, then subtract 8 from both sides to make it x^2-6x=-8

OpenStudy (anonymous):

After that you want to add half of the coefficient of the x term squared to both sides (if that makes any sense). So half of 6 is 3. 3 squared is 9. After this you'll have x^2-6x+9=1

OpenStudy (anonymous):

Now you have a perfect square term on the left which factored will give you (x-3)^2=1

OpenStudy (anonymous):

Now to find the vertex, you look at it in this form as a(x-h)^2+k with vertex (h,k)

OpenStudy (anonymous):

So your vertex will be (3,-1)

OpenStudy (anonymous):

got it thank you

OpenStudy (anonymous):

No problem!